The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Bryan C. Dunne Co-Editors Eva K. Grebel and You-Hua Chu

                                 mcnews@astro.uiuc.edu
                  http://www.astro.uiuc.edu/projects/mcnews/MCNews.php

No. 73 December 10th, 2003

Contents

4 Abstracts of Refereed Papers  2
1 Abstract of a Non-Refereed Paper 4
1 Meeting Announcements  5
3 Job Postings  6
Recent astro-ph Listings  8

Letter From The Editor

Dear Colleagues,

The Magellanic Clouds Working Group is pleased to bring you issue 73 of the Magellanic Clouds Newsletter. In this issue, we present 5 submitted abstracts, an announcement for an upcoming ESO workshop, 3 recent job postings, and a collection of recent astro-ph listings of potential interest to Magellanic Clouds researchers. As always, this month’s issue is available from the MC News Website in a variety of formats (PDF, HTML, PostScript, and L^{A}TEX).

Best Regards, Bryan Dunne Editor, MC News

Abstracts of Refereed Papers

0103-72.6: A New Oxygen-rich Supernova Remnant in the Small Magellanic Cloud

Sangwook Park(1), John P. Hughes(2), David N. Burrows(1), Patrick O. Slane(3), John A. Nousek(1), and Gordon P. Garmire(1)

(1) Penn State University

(2) Rutgers University

(3) Center for Astrophysics

103-72.6, the second brightest X-ray supernova remnant (SNR) in the Small Magellanic Cloud (SMC), has been observed with the Chandra X-Ray Observatory. Our Chandra observation unambiguously resolves the X-ray emission into a nearly complete, remarkably circular shell surrounding bright clumpy emission in the center of the remnant. The observed X-ray spectrum for the central region is evidently dominated by emission from reverse-shockheated metal-rich ejecta. Elemental abundances in this ejecta material are particularly enhanced in oxygen and neon, while less prominent in the heavier elements Si, S, and Fe. We thus propose that 0103-72.6 is a new ”oxygen-rich” SNR, making it only the second member of the class in the SMC. The outer shell is the limb-brightened soft X-ray emission from the swept-up SMC interstellar medium. The presence of O-rich ejecta and the SNR’s location within an H II region attest to a massive star core-collapse origin for 0103-72.6. The elemental abundance ratios derived from the ejecta suggest an 18 Msun progenitor star.

Comments: Published on the ApJ Letters (Dec. 1, 2003), ApJ, 598, L95. WWW: http://www.journals.uchicago.edu/ApJ/journal/issues/ApJL/v598n2/17768/17768.html e-mail: park@astro.psu.edu

Characteristics and classification of A-type supergiants in the Small Magellanic Cloud

C. J. Evans (1), I. D. Howarth(2)

(1) Isaac Newton Group of Telescopes, Santa Cruz de La Palma, Canary Islands, Spain

(2) Department of Physics and Astronomy, University College London, UK

We address the relationship between spectral type and physical properties for A-type supergiants in the SMC. We first construct a self-consistent classification scheme for A supergiants, employing the calcium K to Hϵ line ratio as a temperature-sequence discriminant. Following the precepts of the ‘MK process’, the same morphological criteria are applied to Galactic and SMC spectra with the understanding there may not be a correspondence in physical properties between spectral counterparts in different environments. We then discuss the temperature scale, concluding that A supergiants in the SMC are systematically cooler than their Galactic counterparts at the same spectral type, by up to ∼10%. Considering the relative line strengths of Hγ and the CH G-band we extend our study to F and early G-type supergiants, for which similar effects are found. We note the implications for analyses of extragalactic luminous supergiants, for the flux-weighted gravity-luminosity relationship and for population synthesis studies in unresolved stellar systems.

Comments: MNRAS, 345, 1223 WWW: http://arxiv.org/abs/astro-ph/0308125 e-mail: cje@ing.iac.es

The Star Formation History of the Small Magellanic Cloud

Jason Harris and Dennis Zaritsky

Steward Observatory

We present the spatially-resolved star formation and chemical enrichment history of the Small Magellanic Cloud (SMC) across the entire central 4^{∘} × 4.5^{∘} area of the main body, based on UBV I photometry from our Magellanic Clouds Photometric Survey. We find that 1) approximately 50% of the stars that ever formed in the SMC formed prior to 8.4 Gyr ago (z > 1.2 for WMAP cosmology), 2) the SMC formed relatively few stars between 8.4 and 3 Gyr ago, 3) there was a rise in the mean star formation rate during the most recent 3 Gyr punctuated by “bursts” at ages of 2.5, 0.4, and 0.06 Gyr,

4) the bursts at 2.5 and 0.4 Gyr are temporally coincident with past perigalactic passages of the SMC with the Milky Way, 5) there is preliminary evidence for a large-scale annular structure in the 2.5 Gyr burst, and 6) the chemical enrichment history derived from our analysis is in agreement with the age-metallicity relation of the SMC’s star clusters. Consistent interpretation of the data required an ad hoc correction of 0.1–0.2 mag to the B-V colors of 25% of the stars; the cause of this anomaly is unknown, but we show that it does not strongly influence our results.

Comments: Accepted for publication in AJ WWW: http://marvin.as.arizona.edu/∼jharris/papers/smcsfh.tar.gz e-mail: jharris@as.arizona.edu

Quantifying the Drivers of Star Formation on Galactic Scales. I. The Small Magellanic Cloud

Dennis Zaritsky and Jason Harris

Steward Observatory

We use the star formation history of the Small Magellanic Cloud (SMC) to place quantitative limits on the effect of tidal interactions and gas infall on the star formation and chemical enrichment history of the SMC. The coincident timing of two recent (< 4 Gyr) increases in the star formation rate and SMC/Milky Way(MW) pericenter passages suggests that global star formation in the SMC is driven at least in part by tidal forces due to the MW. The Large Magellanic Cloud (LMC) is the other potential driver of star formation, but is only near the SMC during the most recent burst. The poorly constrained LMC-SMC orbit is our principal uncertainty. To explore the correspondence between bursts and MW pericenter passages further, we model star formation in the SMC using a combination of continuous and tidally-triggered star formation. The behavior of the tidally-triggered mode is a strong inverse function of the SMC-MW separation (preferred behavior ∼ r^{−5}, resulting in a factor of ∼ 100 difference in the rate of tidally-triggered star formation at pericenter and apocenter). Despite the success of these closed-box evolutionary models in reproducing the recent SMC star formation history and current chemical abundance, they have some systematic shortcomings that are remedied by postulating that a sizable infall event (∼ 50% of the total gas mass) occurred ∼ 4 Gyr ago. Regardless of whether this infall event is included, the fraction of stars in the SMC that formed via a tidally triggered mode is > 10% and could be as large as 70%.

Comments: Accepted for publication in AJ WWW: http://xxx.lanl.gov/abs/astro-ph/0312127 e-mail: dzaritsky@as.arizona.edu

Abstracts of Non-Refereed Papers

Monitoring the Evolution of SNR1987A with Chandra

Sangwook Park (1), David N. Burrows (1), Svetozar A. Zhekov (2), Dick McCray (3), and Gordon P. Garmire (1)

(1) Penn State (2) Space Research Institute, Bulgaria (3) Colorado

We report on the results of our monitoring program of SNR 1987A with the Chandra X-Ray Observatory. As of December 2002 we have performed a total of seven observations of SNR 1987A which allows us to monitor the details of the evolution process in the earliest stage of the SNR. The high resolution images from the latest data reveal developments of X-ray bright spots in the north and the west portions as well as in the south-east of the SNR. The observed soft X-ray flux is increasing more rapidly than ever and the latest flux ( 6 x 10**-13 ergs s-1 cm-2) is four times brighter than three years ago when this monitoring began. These results indicate that the shock front has reached the main body of the inner circumstellar ring and that SNR 1987A will be a complete ring with a dramatic brightening in coming years. We present the morphological and spectral evolution of the X-ray remnant of SN 1987A for the last three years.

Comments: To appear in IAU Symposium, 2003, 218E, 154P. WWW: ADS: http://adsabs.harvard.edu/cgi-bin/nph-bib query?bibcode=2003IAUS..218E.154P&db key=AST&high=3be1c07aca23461 e-mail: park@astro.psu.edu

Meeting Announcements

Planetary Nebulae Beyond the Milky Way: An ESO Workshop

May 19–21, 2004 Garching bei München, Germany

We would like to draw to your attention this three day workshop and the first full meeting on planetary nebulae beyond the Milky Way. It is aimed to cover all the current topics in extra-galactic planetary nebula (PN) research from surveys, PN as distance indicators, probes of stellar evolution in different environments, PN as probes of galaxy evolution, use of PN for studying galaxy kinematics and PN in the intra-cluster environment. It will feature invited and contributed talks, posters and discussion sessions.

The format of the meeting will consist of invited reviews and contributed talks in the key areas of extra-galactic PN research. Central topics will be (with confirmed invited speakers):

∘ Historical overview (M. Barlow) ∘ Surveys for extra-galactic PN (R. Corradi, R. Mendez, J. Feldmeier) ∘ The PN luminosity function (R. Ciadullo) ∘ PN in the Magellanic Clouds (G. Jacoby, L. Stanghellini, E. Villaver) ∘ Observational dynamics of PN in galaxies (N. Douglas, E. Peng) ∘ Modelling dark matter with PN dynamical data (O. Gerhard) ∘ Status of abundance determinations in PN (X.-W. Liu) ∘ Dependence of PN evolution on environment (C. Charbonnel) ∘ PN as probes of chemical evolution and star formation history (M. Richer) ∘ Extra-galactic PN and stellar populations ∘ Observations of PN in intra-cluster environment of galaxy clusters (M. Arnaboldi) ∘ Workshop Summary (H. Ford)

Poster presentations are also anticipated and there will be scheduled discussion sessions. The proceedings will be published by Springer in the ESO Astrophysics Symposia series.

Deadlines ∘ 31 January 2004 - First Registration ∘ 15 April 2004 - Final Registration

For more information, contact the meeting organizers at: E-mail: pnconf04@eso.org WWW: http://www.eso.org/gen-fac/meetings/extgalpn04/

Job Postings

Postdoctoral Research Position at Johns Hopkins University

A post-doctoral research position is available to work with Dr. Luciana Bianchi at The Johns Hopkins University, Center for Astrophysical Sciences, Baltimore, MD, USA. The research program involves modeling spectra of hot massive stars in Local Group galaxies with state of the art non-LTE codes. The program includes reduction and analysis of HST-STIS (UV and visual) spectra, groundbased (VLT) and FUSE far-UV spectra of early type stars (and CSPN) in the Local Group. Experience with codes for synthetic spectra calculations is required. Experience in reducing spectroscopic data is a desirable plus.

The candidate should have a PhD in Astronomy or closely-related degree. The position is available immediately (one year, renewable) and will be filled as soon as a suitable candidate is found. Applications including curriculum vitae, statement of interest and relevant work experience, publications, and three letters of reference should be sent to:

Send application materials to Dr. Luciana Bianchi: Address:

The Johns Hopkins University

239 Bloomberg Center for Physics and Astronomy

3400 N. Charles Street

Baltimore, MD 21218-2695

USA E-mail: bianchi@pha.jhu.edu WWW: http://dolomiti.pha.jhu.edu The Johns Hopkins University encourages applications from women and minority candidates. AAE/EOE

Postdoctoral Research Associate in Planetary Nebulae - AURA/NOAO

Postdoctoral Research Associate in Planetary Nebulae The National Optical Astronomical Observatory (NOAO) invites applications for a postdoctoral research associate to work with Letizia Stanghellini on Galactic and Magellanic Cloud planetary nebulae (PNs). Applicants with observational and theoretical background in the evolution of low- and intermediate-mass stars and the physics of diffuse nebulae are particularly encouraged to apply. The postdoc will work on the analysis and the interpretation of space-based (HST) and ground-based data of Galactic and Magellanic Cloud PNs and their central stars. The specific responsibilities of the postdoc will depend on her/his background and interests.

Applicants should have obtained, by the starting date, a Ph.D. in astronomy or a related field. The appointment is for 1 year, with possible extension. Interested applicants are asked to submit by January 15, 2004, a CV, list of publications, and a statement of current and future research interests electronically or by mail to the address shown below. Please also arrange to have three letters of reference sent electronically or by mail to the same address.

Application deadline: January 15th, 2004

Send application materials to:

Address:

Post Doctoral Research Associate in Planetary Nebulae

AURA/NOAO

PO Box 26732

Tucson, AZ 85726-6732

USA Phone: +1-520-318-8000 FAX: +1-520-318-8560 E-mail submissions: hrnoao@noao.edu

E-mail inquiries: lstanghe@stsci.edu NOAO is an equal opportunity and affirmative action employer. We value and foster a diverse research environment. Women and underrepresented minorities are strongly encouraged to apply.

FUSE Research Position-JOHNS HOPKINS UNIVERSITY

The Far Ultraviolet Spectroscopic Explorer (FUSE) group at the Johns Hopkins University is seeking an astronomer interested in analyzing and interpreting scientific data from the FUSE satellite. FUSE is a high resolution ultraviolet spectrograph designed to study the interstellar and intergalactic medium, active galactic nuclei, and stellar and solar system phenomena. The reach of the telescope and data quality are unprecedented in this spectral region. Applicants should have a Ph.D. in astronomy or physics and have demonstrated scientific accomplishments. Previous experience in observational studies of the interstellar medium and astronomical spectroscopy is highly desirable. Applicants will be considered primarily at the Post Doctoral Associate level. Appointment at other levels, (Assistant Research Scientist and Associate Research Scientist) is possible, depending on background and experience. The appointee will have access to the Apache Point Observatory and the Sloan Digital Sky Survey. Further information on the FUSE Project may be obtained at http://fuse.pha.jhu.edu.

Applicants should send a resume and the names of three references to Ms. Lauren Fowler, at the address below as soon as possible. Review of applications will begin 1 February 2004 and will continue until the position is filled.

Application deadline: February 1st, 2004

Send application materials to Ms. Lauren Fowler: Address:

The Johns Hopkins University

151 Bloomberg Center for Physics and Astronomy

3400 N. Charles Street

Baltimore, MD 21218

USA Phone: +1-410-516-2372 FAX: +1-410-516-5494 E-mail: lfowler@pha.jhu.edu WWW: http://fuse.pha.jhu.edu The Johns Hopkins University encourages applications from women and minority candidates. AAE/EOE

Recent astro-ph Listings

astro-ph/0312237:
Title: Red variables in the OGLE-II database. II. Comparison of the Large and
Small Magellanic Clouds
Authors: L.L. Kiss, T. Bedding
Comments: 6 pages, 4 figures, accepted for publication in MNRAS (Pink Pages)
astro-ph/0312233:
Title: Understanding B-type Supergiants in the Low Metallicity Environment of
the SMC
Authors: C. Trundle (1,2), D.J. Lennon (1), J. Puls (3), P.L.Dufton (2) ((1)The
Isaac Newton Group of Telescopes, La Pama, Spain (2)Queen's University of
Belfast, Belfast, N. Ireland (3)Universitats-Sternwarte Munchen, Germany)
Comments: 19 pages, 6 tables, 10 figures, accepted for publication in Astronomy
and Astrophysics (5/12/2003)
astro-ph/0312223:
Title: A New Look at the Kinematics of Neutral Hydrogen in the Small
Magellanic Cloud
Authors: S. Stanimirovic (1), L. Staveley-Smith (2), P. Jones (3) ((1) UC
Berkeley, (2) ATNF)
Comments: To appear in ApJ, March 20 2004, 11 figures
astro-ph/0312127:
Title: Quantifying the Drivers of Star Formation on Galactic Scales. I. The Small
Magellanic Cloud
Authors: Dennis Zaritsky, Jason Harris (Steward Obs., University of Arizona)
Comments: Accepted for publication in ApJ
astro-ph/0312112:
Title: Formation of star clusters in the LMC and SMC. I. Preliminary results on
cluster formation from colliding gas clouds
Authors: Kenji Bekki, Michael A. Beasley, Duncan A. Forbes, Warrick J. Couch
Comments: 22 pages, 5 figures (two color: fig2 and fig5), accepted in ApJ
astro-ph/0312100:
Title: The Star Formation History of the Small Magellanic Cloud
Authors: J. Harris, D. Zaritsky
Comments: 35 pages, 12 Figures; accepted for publication in AJ. Postscript
version with high-resolution Figures and complete machine-readable Table 2
available at this http URL
astro-ph/0312079:
Title: Variable stars in the Magellanic Clouds: II. The data and infrared
properties
Authors: Yoshifusa Ita (1), Toshihiko Tanabe (1), Noriyuki Matsunaga (1),
Yasushi Nakajima (2), Chie Nagashima (2), Takahiro Nagayama (3), Daisuke
Kato (3), Mikio Kurita (3), Tetsuya Nagata (3), Shuji Sato (3), Motohide
Tamura (2), Hidehiko Nakaya (4), Yoshikazu Nakada (1,5) ((1)University of
Tokyo (2)NAOJ (3)Nagoya University (4)Subaru, NAOJ (5)Kiso Observatory)
Comments: References added. Submitted to MNRAS. 6 pages, 7 figures. High
resolution version (about 860kb) is available at: this http URL

astro-ph/0312051:
Title: Massive Field Stars and the Stellar Clustering Law
Authors: M. S. Oey (Lowell), N. L. King (Lowell), J. Wm. Parker (SWRI)
Comments: Accepted by AJ. 13 pages, 9 figures in 11 files, uses emulateapj.sty
astro-ph/0312034:
Title: High Mass X-ray Binary Pulsars in the Small Magellanic Cloud
Authors: W.A. Majid (1), R.C. Lamb (2), D.J. Macomb (3) ((1) Jet Propulsion
Laboratory, California Institute of Technology, (2) Space Radiation Laboratory,
California Institute of Technology, (3) Physics Department, Boise State
University)
Comments: submitted to The Astrophysical Journal
astro-ph/0311548:
Title: Resolution of the MACHO-LMC-5 Puzzle: The Jerk-Parallax Microlens
Degeneracy
Authors: Andrew Gould (Ohio State)
Comments: 17 Pages, Submitted to ApJ
astro-ph/0311477:
Title: Mass Segregation in Young Magellanic Clouds Star Clusters: Four
Clusters observed with HST
Authors: D. Gouliermis, S. C. Keller, M. Kontizas, E. Kontizas, I.
Bellas-Velidis
Comments: A&A Accepted, 20 pages, 9 Figures, Version with language errors
and typos corrected
astro-ph/0311322:
Title: Were the LMC Globular Clusters Formed in a Disk?
Authors: Sidney van den Bergh (Dominion Astrophysical Observatory,
Herzberg Institute of Astrophysics, National Research Council of Canada)
Comments: 6 pages, 1 figure, Publication in the February 2004 issue of the
Astronomical Journal
astro-ph/0311157:
Title: Nonradial Pulsations in Classical Cepheids of the Magellanic Clouds
Authors: Pawel Moskalik, Zbigniew Kolaczkowski, Tomasz Mizerski
Comments: 4 pages, 2 figures, to appear in "Variable Stars in the Local Group",
ASP Conference Series, eds. D. W. Kurtz & K. Pollard
astro-ph/0311092:
Title: X-ray observations of Be/X-ray binaries in the SMC
Authors: F. Haberl, W. Pietsch
Comments: 10 pages, 7 figures, accepted for publication in A&A
astro-ph/0310860:
Title: Evidence for a fundamental stellar upper mass limit from clustered star
formation
Authors: Carsten Weidner, Pavel Kroupa (University if Kiel)
Comments: MNRAS, accepted, 6 pages
astro-ph/0310807:
Title: Magellanic Cloud planetary nebulae as probes of stellar evolution and
populations
Authors: Letizia Stanghellini
Comments: to be published in the proceedings of "The Local Group as an
Astrophysical Laboratory", May Symposium held in May 2003, STScI

astro-ph/0310803:
Title: Magellanic Cloud PNs as probes of stellar evolution and populations
Authors: Letizia Stanghellini
Comments: to appear in "Asymmetrical Planetary Nebulae III" editors M.
Meixner, J. Kastner, N. Soker, & B. Balick (ASP Conf. Series)
astro-ph/0310720:
Title: Once again about the origin of the system of the giant stellar arcs in the
Large Magellanic Cloud
Authors: Yu.N.Efremov
astro-ph/0310713:
Title: Thermal and Non-thermal X-Rays from the LMC Super Bubble 30 Dor C
Authors: Aya Bamba, Masaru Ueno, Hiroshi Nakajima, Katsuji Koyama (Kyoto
Univ.)
Comments: 11 pages, 6 figures, Accepted for publication in ApJ, the paper with
full resolution images in this http URL
astro-ph/0310673:
Title: A Review of the Distance and Structure of the Large Magellanic Cloud
Authors: David R. Alves
Comments: ver. 2, updated refs., 10 + 2 pages, now incl. Appendix w/ long-form
abstract for "Highlights of Astronomy, Vol. 13," ed. O. Engvold after review
paper for Proc. of JD 13 on "Extragalactic Binaries," eds. A. Gimenez and I.
Ribas, invited talk from IAU XXV General Assembly
astro-ph/0310582:
Title: Beta Cephei and SPB stars in the Large Magellanic Cloud
Authors: Z. Kolaczkowski, A. Pigulski, I. Soszynski, A. Udalski, M. Szymanski,
M. Kubiak, K. Zebrun, G. Pietrzynski, P.R. Wozniak, O. Szewczyk, L.
Wyrzykowski
Comments: 4 pages, 3 figures; to appear in the proceedings of the IAU
Colloquium 193 `Variable Stars in the Local Group', eds. Don Kurtz and Karen
Pollard
astro-ph/0310562:
Title: First results from a Chandra survey of the 'Bar' region of the SMC
Authors: A. Zezas (1), J. C. McDowell (1), D. Hadzidimitriou (2), V. Kalogera
(3), G. Fabbiano (1), P. Taylor (1,4) ((1) CfA, (2) University of Crete, (3)
Northwestern University, (4) Boston College)
Comments: Poster presented at the 2003 STScI Spring Symposium "The Local
Group as an Astrophysical Laboratory"
astro-ph/0310545:
Title: Distance Scale, Variable Stars and Stellar Populations in Local Group
Galaxies
Authors: G. Clementini, et al. (INAF - Bologna Observatory)
Comments: 4 pages, 4 figures, Proceedings of the IAU Colloq. 193, "Variable
Stars in the Local Group", ed. D. W. Kurtz & K. R. Pollard., ASP Conf. Ser., in
press

astro-ph/0310459:
Title: Comprehensive modelling of the planetary nebula LMC-SMP 61 and its
[WC]-type central star
Authors: G. Stasi\'nska (1), G. Grfener (2), M. Pea (3), W.-R. Hamann (4),
L. Koesterke (4), R. Szczerba (5) ((1) LUTH, Observatoire de Meudon, France
(2) Institut fr Physik, Astrophysik, Universitt Potsdam, Germany (3) Instituto
de Astronom{\'\i}a, UNAM, Mxico (4) Laboratory for Astronomy and Solar
Physics, NASA GSFC, USA (5) N. Copernicus Astronomical Center, Toru n,
Poland)
Comments: 16 pages, 10 figures, Astronomy and Astrophysics in press
astro-ph/0310405:
Title: Do we really observe a bow shock in N157B...?
Authors: Eric van der Swaluw
Comments: 6 pages, 3 figures, to appear in "High Energy Studies of Supernova
Remnants and Neutron Stars" (COSPAR 2002), Advances in Space Research, in
press
astro-ph/0310349:
Title: LMC Bar: Evidence for a warped bar
Authors: Annapurni Subramaniam
Comments: 10 pages with 4 figures. ApJ Letters (in press)
astro-ph/0310313:
Title: An Intriguing X-ray Arc Surrounding the X-ray Source RX
J053335-6854.9 toward the Large Magellanic Cloud
Authors: J.D. Lowry, Y.-H. Chu, M.A. Guerrero, R.A. Gruendl (University of
Illinois), S.L. Snowden (NASA/GSFC), R.C. Smith (CTIO/NOAO)
Comments: 13 pages, 5 figures, accepted for publication in the Astronomical
Journal, 2004 Jan issue
astro-ph/0310281:
Title: The MACHO Project Large Magellanic Cloud Variable Star Inventory.
XIII. Fourier Parameters for the First Overtone RR Lyrae Variables and the
LMC Distance
Authors: C. Alcock, D. R. Alves, T. S. Axelrod, A. C. Becker, D. P. Bennett, C.
M. Clement, K. H. Cook, A. J. Drake, K. C. Freeman, M. Geha, K. Griest, M. J.
Lehner, S. L. Marshall, D. Minniti, A. Muzzin, C. A. Nelson, B. A. Peterson, P.
Popowski, P. J. Quinn, A. W. Rodgers, J. F. Rowe, W. Sutherland, T. Vandehei,
D. L. Welch
Comments: 51 pages, 17 figures, 12 tables, accepted to AJ
astro-ph/0310217:
Title: Determining the Metallicity of Cepheid Stars in the SMC, LMC and the
Galaxy
Authors: M. Mottini (1), F. Primas (1), M. Romaniello (1), M. Groenewegen
(2), ((1) ESO-Garching, (2) Instituut voor Sterrenkunde, Leuven)
Comments: 4 pages, 1 figure, to appear in "Variable Stars in the Local Group",
IAU Colloquium 193 proceedings, ASP Conference Series, D.W. Kurtz & K.
Pollard eds
astro-ph/0310199:
Title: Coupling between satellite dwarfs and the Milky Way warp
Authors: Jeremy Bailin (1), Matthias Steinmetz (1 and 2) ((1) Steward
Observatory, University of Arizona, (2) Astrophysikalisches Institut Potsdam)
Comments: 5 pages, 2 figures, to appear in the proceedings of the conference
"Satellites and Tidal Streams"